Diagraph of Metaphysic or Ontology

Jacob Lorhard, trans. by Sara L. Uckelman^{∗}

The parts of metaphysic (which is GREEK, because from natural man by the light of reason without any conception of matter it is intelligible.) are two; Either

\begin{cases}Universal, which consists of intelligi- bles and beings. & \begin{cases}By most general distributions. How- ever it should be noted of an Intelligible. & \begin{cases}λ�γος: An Intelligible is said to be anything, which is perceived and comprehended by the intellect. \\ Distribution: An Intelligible is either & \begin{cases}Nothing: This is simply not something. \\ Something: What- ever is simply not nothing. It is either & \begin{cases}Positive, because it fixes or affirms something. It is either & \begin{cases}Essence. See A. \\ Being. See B. \end{cases} \\ Negative. See RRR. \end{cases} \end{cases} \end{cases} \\ By most common attributis. See C. \end{cases} \\ Particular. See EE. \end{cases}

∗Many thanks to Brian M. Scott and Ursula Whitcher for help with one or two tricky bits, and Samson de Jager and Joel Uckelman for putting in nearly as much time making the L^{A}TEXpackage doing this display as I did translating.

A. Of this it should be observed \begin{cases}λ�γος: Essence is something productive, by which a Being is what it is. This foreign [word] is called by the Scholastics by the word Entity or quidditas. \\ Distribution. For example essence is either \{Real; which is an actu- ality of a real being, because it is of such a kind. Imaginary; which is an actuality of an rational being, be- cause it is of such a kind.\} Both are & \{ \begin{matrix}1 & \begin{cases}Simple, which belongs to a simple being. \\ Composite, which belongs to a composite being. \end{cases} \\ 2 & \begin{cases}Absolute, which is of an absolute being. \\ Respective, which is of a respective being. \end{cases} \\ 3 & \begin{cases}Universal, which is of a universal being. It is ei- ther & \begin{cases}Generic, in so far as a being, by its own genus, has [things] common with other species. \\ Specific, through which a being is placed in a certain species. \end{cases} \\ Singular, which belongs to any singular and individual being. \end{cases} \end{matrix} \end{cases}

2